OpenStudy (anonymous):

How do you find the 2nd derivative of xy+e^y=e?

OpenStudy (yttrium):

Derive the equation twice. So far, where are you? @Jmen7164

OpenStudy (anonymous):

I got the first derivative as y/(e-xy-x), which I am positive is incorrect.

OpenStudy (yttrium):

First derivative (using implicit differentiation) \[xy+e^y = e\] \[(y+xy')+e^y y' = 0\] \[y' = \frac{ -y }{ x+e^y }\] Second derivative: Derive the first derivative. \[y'' = \frac{ (x+e^y)(-y')-(-y)(1+e^yy') }{ (x+e^y)^2 }\] Therefore, \[y''=\frac{ (x+e^y)(-(\frac{ -y }{ x+e^y })+ y(1+e^y(\frac{ -y }{ x+e^y }) }{ (x+e^y)^2 }\] And upon simplifying we'll have. \[y''=\frac{ 2y(x+e^y)-y^2e^y }{ (x+e^y)^2 }\]

OpenStudy (yttrium):

@Jmen7164 can you follow?

OpenStudy (anonymous):

I guess I am just confused because the back of the book also says the answer is 1/e^2 and I cannot simply an answer to get that.

OpenStudy (anonymous):

What class are you in? And are you sure you're not evaluating it at some point?

OpenStudy (anonymous):

I would be evalutating it at 0, but wouldn't that still leave me with -y^2e^y/(e2y)

OpenStudy (anonymous):

and is it dx/dy or dx/dy?

OpenStudy (anonymous):

They're wanting you to make it x = (e-e^y)/y then finding the second derivative of that, yah?

OpenStudy (yttrium):

Oh my. I just lost my efforts, if that's the case :(

OpenStudy (anonymous):

You're definitely more right than me...